I am stuck on correct answer for:
(1 - 0.7) - 7.47 =
I believe the answer is -7.17, What is correct answer to this?
I am not using any specific book. I am just practicing...

Answer by jenrobrody(19)

You are correct.
(1 - 0.7) - 7.47 =
(1.0 + (-0.7)) - 7.47 =
(0.3) + (-7.47) =
-7.17